OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. x – 3y = –8 x – 3y = –9 3x + y = 9 3x – 3y = –8

OpenStudy (anonymous):

First you need to solve for the slope of the ordered pairs using the euqation: y2 - y1/ x2-x1 = M Then solve for the equation (y - y1) = m (x - x1) You should get an equation and you could solve with that to figure which option it would be =)

OpenStudy (anonymous):

the equation comes to: y-3=1/3(x-1) the first step to solve you subtract 3 but then im stuck